OpenStudy (anonymous):

solve the system by graphing. -2x=2y-4 2x-y=-5 leaseeee help! tell me if you need the answer choices

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

also some help with some other questions would be much appreciated?

OpenStudy (anonymous):

y=-x+4 y=2x+5 Slope of line 1 = -2, 4 is y intercept Slope of line 2= 2, 5 is y intercept. Intersection of lines is solution.

OpenStudy (anonymous):

so how would that look on a graph? what would the coordinates be?

OpenStudy (anonymous):

i know their intersecting lines, just unsure of the coordinates..

OpenStudy (anonymous):

Either read it off the graph, or if they're in between points, y1=y2 -x+4=2x+5 Solve for x, plug into equation for y coord.

OpenStudy (dumbcow):

@nukka , typo first line should be "y = -x +2"

OpenStudy (anonymous):

y=-x+2 y=2x+5 Slope of line 1 = -1, 2 is y intercept Slope of line 2= 2, 5 is y intercept. Either read it off the graph, or if they're in between points, y1=y2 -x+2=2x+5 Solve for x, plug into equation for y coord.

OpenStudy (anonymous):

-1,3 3,-1 -3,-1 -1,-3 these are the answer choices?

OpenStudy (anonymous):

2x+x=2-5 x=-1,y=3

OpenStudy (anonymous):

from what was said before mean that the first coordinate would be negative and the second positive?

OpenStudy (anonymous):

oh okay thank you!